I need help with these problems.
Solve the unknown 3/5=x/15
complete the statement 25cl= mil
sides of a square =13 5/6 find the perimeter

Answer by rfer(16322)

3/5=x/15
3/5*15=x
x=45/5
x=9
----------------
250 mil
------------
13 5/6*4=x
x=83/6*4
x=332/6
x=55 1/3